15 basketball players are to be drafted by 3 professional teams (A,B,C). Each team will draft 5 players. In how many ways can this be done?
Answer by Edwin McCravy(20054)

15 basketball players are to be drafted by 3 professional teams (A,B,C). Each team
will draft 5 players. In how many ways can this be done?
We can choose 5 players for team A in any of 15 ways. That's 15C5 = 3003 ways.

We can then choose 5 players for team B in any of the remaining 10 ways. That's
10C5 = 252 ways.

The remaining 5 players will automatically go to team C.  That's 5C5 = 1 way.

Answer 15C5*10C5*5C5 = 3003*252*1 = 756756

Edwin